APPLICATION OF THE DIABATIC CORRELATION METHOD TO LABELING SPECTRA OF MOLECULES CONTAINING AN INTERNAL ROTOR

J. ORTIGOSO, Instituto de Estructura de la Materia, Consejo Superior de Investigaciones Cientificas, Serrano 121, 28006 Madrid, Spain; \underline{J. T. HOUGEN}, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD. 20899, USA.

Labeling problems for rotational energy levels of asymmetric-top molecules containing a noncoaxial internal rotor are investigated by using Rose and Kellman’s diabatic correlation methods. We find that v_{t} = 2 and v_{t} = 3 rotation-torsion eigenstates for acetaldehyde, which lie just below and just above the barrier to internal rotation, can be unambiguously and meaningfully labeled by nominal quantum numbers K_{a} corresponding to eigenstates of a suitably chosen zeroth-order symmetric-top Hamiltonian. Such a K_{a} labeling is shown to be superior to that based on eigenvector composition for this near-prolate asymmetric rotor. The K_{a} assignments from the diabatic correlation method agree with assignments made using the criterion, based on spectroscopic intuition, of smooth variation of the B value along a series of levels of given K_{a} and increasing J. As a second example, the v_{t} = 0 state of acetamide lies near the top of the very low barrier in this near oblate top molecule. For these torsion-rotation states also, meaningful K_{a} labels can be determined from the diabatic correlation method. A discussion of the method and of the results obtained for these two example molecules will be presented.